OpenStudy (diamondboy):

I use a thermometer graduated in 1/5 degree Celsius to measure outside air temperature. Measured to the nearest 1/5 degree, yesterday's temperature was 22.4 degrees Celsius and today's is 24.8 degrees Celsius. What is the relative uncertainty in the temperature difference between yesterday and today?

OpenStudy (diamondboy):

I got 16.7 percent

OpenStudy (diamondboy):

@Michele_Laino

OpenStudy (diamondboy):

@freckles

OpenStudy (diamondboy):

@Preetha

OpenStudy (diamondboy):

@SolomonZelman

OpenStudy (diamondboy):

can you help @jdoe0001

OpenStudy (diamondboy):

I would medal

OpenStudy (diamondboy):

@phi

OpenStudy (phi):

I am thinking the temp difference is 2.4 deg the uncertainty is 0.2 on the first measurement and 0.2 on the 2nd so the relative uncertainty *might* be 0.4/2.4 however, I think people combine uncertainty to get sqr(2) * 0.2 (rather that 2*0.2) if that is correct (I'm a bit hazy on this), you would get 0.2828/2.4 = 11.8%

OpenStudy (diamondboy):

the answer is actually 7.7 perccent

OpenStudy (phi):

If you know how they got that number, please post the math

OpenStudy (diamondboy):

ok. I would figure it out

jhonyy9 (jhonyy9):

24,8-22,4=2,4 22,4 ---- 100% 2,4 ----- x% ----------------- x = 2,4*100/22,4 = 10,71 %

OpenStudy (diamondboy):

wrong

OpenStudy (diamondboy):

I am sorry